OpenStudy (anonymous):

Which expression is equivalent to -(x^-2)^-4 a) -1/x^8 b) 1/x^8 c) -x^8 d) x^8

OpenStudy (anonymous):

c

mathslover (mathslover):

\[-(x^{-2})^{-4}\] \[-(1/x^2)^{-4}\] \[-(1^{-4})/(x^2*{-4})\] \[-1/{x^{-8}}\] \[x^8*-1\] \[-x^8\] c is the answer

OpenStudy (anonymous):

is (b)

OpenStudy (mertsj):

When you raise to a power, you multiply the exponents. \[-(x ^{-2})^{-4}=-(x^8)=-x^8\]

OpenStudy (anonymous):

That's what the law of negative exponents states @Mertsj. It's say that to make it positive you have to flip it.

OpenStudy (anonymous):

@mathslover the result will be like this\[-x^{-8} = -x ^{-8}\]

OpenStudy (anonymous):

so when it will go down the negative sign will change into positive \[1 \over x ^{8}\]

mathslover (mathslover):

@izanagi.wielder i dont think that u r going on right way ... - (x^-8) = -1/(-x^8)

OpenStudy (mertsj):

Stop being ridiculous. The answer is -x^8

OpenStudy (anonymous):

sorry :]

OpenStudy (anonymous):

Oh and I mean to say that's not what the...

OpenStudy (mertsj):

\[-(x ^{-2})^{-4}=-(\frac{1}{x^2})^{-4}=-(x^2)^4=-x^8\]